# Angular momentum of vacuum bubbles in a first-order phase transition

### v1

@@ 4.2.5 Numerical scheme
To determine the transfer functions during the FOPT, we first numerically solve the background evolution equations, Eqs. (4.28)-(4.33), and then work out the perturbation equations, given in Eqs. (4.57)-(4.60) for a given mode $k/\mathcal{H}_{\rm c}$ and for a fluid component (SM or dark sector), up to the percolation time. Since the perturbations are crucial for the calculation of the angular momentum of FV remnants, it is necessary to perform checks on our numerical scheme. We can compare our numerical transfer functions
$U_{Q_{\rm{I}}}\equiv\frac{Q_{\rm I}}{\mathcal{R}_{k}}\,,\quad Q_{\rm I}=\delta_{\rm D},~{}\hat{\psi}_{\rm D},~{}\Phi,~{}\nabla^{2}\gamma\,,$
with the WKB approximation and the SM perturbations, for subhorizon modes.
In Fig. 9 we show plots of $U_{\delta}$ (top row) and $U_{\hat{\psi}}$ (bottom row) as functions of the rescaled elapsed conformal time, for BP-2 and BP-7, as in Fig. 3, where we fix $k/\mathcal{H}_{\rm c}=1000$.
@@
Figure 9: The density (top row) and velocity (bottom row) perturbations for benchmarks BP-2 (column a) and BP-7 (column b) for $k/\mathcal{H}_{\rm c}=10^{3}$.
In all panels, the solid curves correspond to the dark sector perturbation, the dashed curves correspond to the corresponding SM perturbation, and the dotted curves correspond to the WKB approximation. In all cases, we observe excellent agreement between the WKB approximation and the numerical solution. There is a deviation between the two results close to the percolation time, which is at the right-most edge of each panel. This discrepancy between the numerical solution and the WKB approximation can be understood by noting that $c_{s\rm,D}^{2}$ is no longer a slowly varying function close to the percolation time.
@@
Figure 10: The absolute sizes of $U_{\delta},U_{\hat{\psi}}$ at the percolation time, for a range of modes $10^{-3}<k/\mathcal{H}_{\rm c}<(k/\mathcal{H}_{\rm c})_{\rm cut}$, and taking FOPT benchmarks BP-2 (column a) and BP-7 (column b) as in Fig. 3.
Ultimately, in calculating the angular momentum of FV remnants, we need the transfer functions at the percolation time. In Fig. 10, we show the absolute values of the transfer functions, evaluated at the percolation time, for modes within $0.1<k/\mathcal{H}_{\rm c}<k_{\rm cut}$, where $k_{\rm cut}=5/(H_{*}R_{*})$, $H_{*}$ is the Hubble parameter at percolation, and $R_{*}$ is the physical radius of a FV bubble at percolation. We take the physical radius $R_{*}$ of the FV bubble at percolation to be
@@
where $n_{\rm FV}(t_{*})$ is the number density of FV bubbles at the percolation time. Equivalently, this critical radius $R_{*}$ is determined from the condition that no TV bubble nucleates inside, so that
$\Gamma(t_{*})~{}\frac{R_{*}}{v_{\rm w,*}}~{}\frac{4\pi R_{*}^{3}}{3}=1\,.$ (4.84)
It is straightforward to show that
@@
In panels (a) and (b) of Fig. 10, we can see that the superhorizon modes follow the expected scaling behaviors $\delta_{\rm D}\propto(k/\mathcal{H})^{2}$ and $\hat{\psi}_{\rm D}\propto(k/\mathcal{H})^{3}$. In the subhorizon regime $k/\mathcal{H}_{\rm c}\gtrsim 1$, the rapid oscillations of $U_{\delta}$ and $U_{\hat{\psi}}$ are manifestations of the WKB behavior of the solutions in the deep subhorizon regime. For perturbative modes $k/\mathcal{H}_{\rm c}>10^{2}$, we observe an enhancement in the transfer functions for the density perturbations. We can attribute this to the period when $c_{s\rm,D}^{2}$ exhibits a dip around the percolation time, as seen in the bottom row of Fig. 3. This behavior is similar to the main results of [21], where the vanishing sound speed of the plasma leads to an enhancement in the density perturbations. In both benchmarks BP-2 and BP-7, the width of the dip is $\mathcal{H}_{\rm c}(\eta_{f}-\eta_{i})\simeq 4.2\times 10^{-3}$, indicating that those modes with wavenumbers of around at least $k/\mathcal{H}_{\rm c}\simeq\frac{1}{\mathcal{H}_{\rm c}(\eta_{f}-\eta_{i})}\simeq 250$, will have enough resolution to resolve the dip feature in the temporal evolution of the sound speed.
Since the WKB approximation breaks down at the percolation time, we can understand the behavior of the transfer functions in the subhorizon regime by accounting for the sharp decrease in the sound speed at around the percolation time. We solve the ODE,
$\delta_{\rm D}^{\prime\prime}+k^{2}c_{s\rm,D}^{2}\delta_{\rm D}\approx 0\,,$ (4.86)
where we have implicitly assumed that the drag term can be dropped. This is justified provided that
$\left(3c_{s\rm,D}^{2}-6w_{\rm D}+1\right)\left(\frac{k}{\mathcal{H}}\right)c_{s\rm,D}\ll\left(\frac{k}{\mathcal{H}}\right)^{2}c_{s\rm,D}^{2}+9\mathcal{H}^{2}\left(c_{s\rm,D}^{2}-w_{\rm D}\right)+3\mathcal{H}\left(c_{s\rm,D}^{2}\right)^{\prime}\,.$ (4.87)
Then we approximate $c_{s\rm,D}^{2}$ by a piecewise linear function where
$c_{s\rm,D}^{2}\approx\begin{cases}\frac{1}{3},~{}0\leq\tilde{\eta}\leq\tilde{\eta}_{i}\\ ~{}\\ \frac{1}{3}+\left(\frac{c_{s\rm,D*}^{2}-1/3}{\tilde{\eta}_{*}-\tilde{\eta}_{i}}\right)\left(\tilde{\eta}-\tilde{\eta}_{i}\right),~{}\tilde{\eta}_{i}<\tilde{\eta}\leq\tilde{\eta}_{*}\,.\end{cases}$ (4.88)
$\delta=\delta_{\rm c}\cos\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\tilde{\eta}\right)+\sqrt{3}\hat{\psi}_{\rm c}\sin\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}\tilde{\eta}\right)\,,$ (4.89)
where $\delta_{\rm c}$ and $\hat{\psi}_{\rm c}$ are the perturbations evaluated at the critical point, assuming the analytic expression for perturbations evolving in a pure radiation background. On the other hand, in the interval $\tilde{\eta}_{i}<\tilde{\eta}\leq\tilde{\eta}_{*}$, the ODE admits a general solution in terms of Airy functions
$\delta=c_{1}Ai(u)+c_{2}Bi(u),\quad u\equiv\frac{-d_{1}-d_{2}\tilde{\eta}}{(-d_{2})^{2/3}}\,,$ (4.90)
where
@@
$d_{1}=\frac{1}{3}\left(\frac{k}{\mathcal{H}_{\rm c}}\right)^{2}\left[1-3\tilde{\eta}_{i}\left(\frac{c_{s\rm,D*}^{2}-1/3}{\tilde{\eta}_{*}-\tilde{\eta}_{i}}\right)\right]\,,\quad d_{2}=\left(\frac{k}{\mathcal{H}_{\rm c}}\right)^{2}\left(\frac{c_{s\rm,D*}^{2}-1/3}{\tilde{\eta}_{*}-\tilde{\eta}_{i}}\right)\,.$ (4.91)
The coefficients $c_{1}$ and $c_{2}$ are obtained by matching the solutions at $\tilde{\eta}=\tilde{\eta}_{i}$. The density and velocity perturbations at the percolation time will then be
$\delta_{*}=c_{1}Ai(u_{*})+c_{2}Bi(u_{*})\,,\quad\hat{\psi}_{*}=\frac{1}{k/\mathcal{H}_{\rm c}}\left(-d_{2}\right)^{1/3}\left[c_{1}Ai^{\prime}(u_{*})+c_{2}Bi^{\prime}(u_{*})\right]\,,$ (4.92)
where $u_{*}\equiv(-d_{1}-d_{2}\tilde{\eta}_{*})(-d_{2})^{-2/3}$.
We can further understand the approximate transfer functions by performing an asymptotic expansion in powers of $k/\mathcal{H}_{\rm c}$. We find
$\delta_{*}$ $\approx-\frac{4}{(3c_{s\rm,D*}^{2})^{1/4}}\cos\phi_{-}$
@@
$+\frac{1}{k/\mathcal{H}_{\rm c}}\Bigg{[}\left(\frac{5\sqrt{3}(c_{s}^{2})^{\prime}-288\left(c_{s\rm,D*}^{2}\right)^{3/2}+9(c_{s}^{2})^{\prime}\left(c_{s\rm,D*}^{2}\right)^{3/2}}{12~{}3^{3/4}\left(c_{s\rm,D*}^{2}\right)^{7/4}}\right)\sin\phi_{-}$
$-\frac{3}{2}\left(\frac{3}{c_{s\rm,D*}^{2}}\right)^{1/4}(c_{s}^{2})^{\prime}\sin\phi_{+}\Bigg{]}\,,$ (4.93)
and
$\hat{\psi}_{*}$ $\approx\frac{4}{\sqrt{3}}(3c_{s\rm,D*}^{2})^{1/4}\mathcal{R}_{k}\sin\phi_{-}$
$+\frac{\mathcal{R}_{k}}{k/\mathcal{H}_{\rm c}}\Bigg{[}\left(\frac{7\sqrt{3}(c_{s}^{2})^{\prime}+288\left(c_{s\rm,D*}^{2}\right)^{3/2}-9(c_{s}^{2})^{\prime}\left(c_{s\rm,D*}^{2}\right)^{3/2}}{12~{}3^{3/4}\left(c_{s\rm,D*}^{2}\right)^{5/4}}\right)\cos\phi_{-}$
$-\frac{3}{2}\left(3c_{s\rm,D*}^{2}\right)^{1/4}\cos\phi_{+}\Bigg{]}\,.$ (4.94)
Here
ϕ±subscriptitalic-$\phi_{\pm}$$\equiv$ $\frac{1+\tilde{\eta}_{i}}{\sqrt{3}}\frac{k}{\mathcal{H}_{c}}\pm\frac{2}{27(c_{s}^{2})^{\prime}}\left(\sqrt{3}-9\left(c_{s\rm,D*}^{2}\right)^{3/2}\right)\frac{k}{\mathcal{H}_{\rm c}}\,,$ (4.95)
$(c_{s}^{2})^{\prime}$$\equiv$ $-\frac{1/3-c_{s\rm,D*}^{2}}{\tilde{\eta}_{*}-\tilde{\eta}_{i}}\,.$ (4.96)
By comparison, the asymptotic values for $k/\mathcal{H}_{\rm c}\gg 1$ in the pure radiation background case are
@@
$\delta_{*}\rightarrow-4\mathcal{R}_{k}\cos\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}(1+\tilde{\eta}_{*})\right)\,,\quad\hat{\psi}_{*}\rightarrow\frac{4}{\sqrt{3}}\mathcal{R}_{k}\sin\left(\frac{1}{\sqrt{3}}\frac{k}{\mathcal{H}_{\rm c}}(1+\tilde{\eta}_{*})\right)\,.$ (4.97)
We recover the result for radiation when $c_{s\rm,D*}^{2}\rightarrow 1/3$. For $k/\mathcal{H}_{\rm c}\gg 1$, we find that the asymptotic value of $|\delta_{*}|$ is enhanced relative to the pure radiation background by a factor of $(3c_{s\rm,D*}^{2})^{-1/4}$. Similarly, $|\hat{\psi}_{*}|$ will be suppressed by the same factor. These estimates for the enhancement/suppression in the density/velocity transfer functions, for sufficiently large $k/\mathcal{H}_{\rm c}$, are shown as horizontal dashed lines in Fig. 10. We clearly see that the estimates match the results of numerical calculation.
@@
$\tilde{\eta}_{i}=0.1-10^{-3},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=2\times 10^{-3}$ $\tilde{\eta}_{i}=0.1-10^{-4},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=2\times 10^{-3}$

### v2

@@ 4.2.5 Numerical scheme
To determine the transfer functions during the FOPT, we first numerically solve the background evolution equations, Eqs. (4.28)-(4.34), and then work out the perturbation equations, given in Eqs. (4.60)-(4.63) for a given mode $k/\mathcal{H}_{\rm c}$ and for a fluid component (SM or dark sector), up to the percolation time. Since the perturbations are crucial for the calculation of the angular momentum of FV remnants, it is necessary to perform checks on our numerical scheme. In particular we compare our numerical transfer functions for the dark sector density and velocity perturbations, i.e.,
$U_{Q_{\rm{I}}}\equiv\frac{Q_{\rm I}}{\mathcal{R}_{k}}\,,\quad Q_{\rm I}=\delta_{\rm D},\hat{\psi}_{\rm D}\,$,
with (a) the approximate solutions in the subhorizon limit, given by Eqs. (4.84) and (4.85); and (b) the corresponding SM perturbations for modes with $k/\mathcal{H}_{\rm c}\gg 1$.
In Fig. 8 we show plots of $U_{\delta}$ (top row) and $U_{\hat{\psi}}$ (bottom row) as functions of the rescaled elapsed conformal time, for BP-2 and BP-4, as in Fig. 3, where we fix $k/\mathcal{H}_{\rm c}=100$.
@@
Figure 8: The density (top) and velocity (bottom) perturbations for BP-2 (column a) and BP-4 (column b) for $k/\mathcal{H}_{\rm c}=10^{2}$.
In all panels, the solid red curves correspond to the dark sector perturbation, the blue correspond to the corresponding SM perturbation, and the green curves correspond to the approximate solution in the subhorizon limit, which we denote as the Airy function solution. In all cases, we observe excellent agreement between the Airy function approximation, the SM perturbation, and the numerical solution, in the time interval when the sound speed is close to $1/\sqrt{3}$. On the other hand, during the period when the sound speed squared quickly drops, the Airy function approximation captures the behavior of the numerical solution very well.
@@
Figure 9: The absolute sizes of $U_{\delta},U_{\hat{\psi}}$ at the percolation time, for $0.1<k/\mathcal{H}_{\rm c}<k_{\rm max}/\mathcal{H}_{\rm c}$, and BP-2 (column a) and BP-4 (column b). In each panel we show the dark sector perturbation (red), SM sector perturbation (blue), and Airy function approximation (green) to the perturbation.
Ultimately, in calculating the angular momentum of FV remnants, we need the transfer functions at the percolation time. From a practical standpoint, only a finite number of modes can be calculated numerically. We have chosen to solve the perturbation equations up to the percolation time, for $N=1000$ modes between $10^{-3}<k/\mathcal{H}_{\rm c}<k_{\rm max}/\mathcal{H}_{\rm c}$, where $k_{\rm max}\equiv 5/(H_{*}R_{*})$, $H_{*}$ is the Hubble parameter at percolation, and $R_{*}$ is the physical radius of a FV bubble at percolation. In Fig. 9, we show the absolute values of the transfer functions, evaluated at the percolation time, for modes within $0.1<k/\mathcal{H}_{\rm c}<k_{\rm max}/\mathcal{H}_{\rm c}$. We take the physical radius $R_{*}$ of the FV bubble at percolation to be
@@
where $n_{\rm FV}(t_{*})$ is the number density of FV bubbles at the percolation time. Equivalently, this critical radius $R_{*}$ is determined from the condition that no TV bubble nucleates inside,
$\Gamma(t_{*})~\frac{R_{*}}{v_{\rm w,*}}~\frac{4\pi R_{*}^{3}}{3}=1\,$, (4.91)
i.e.,
@@
From panels (a) and (b) of Fig. 9, we can see that the superhorizon modes follow the expected scaling behaviors $\delta_{\rm D}\propto(k/\mathcal{H})^{2}$ and $\hat{\psi}_{\rm D}\propto(k/\mathcal{H})^{3}$. In the subhorizon regime $k/\mathcal{H}_{\rm c}\gtrsim 1$, the rapid oscillations of $U_{\delta}$ and $U_{\hat{\psi}}$ are manifestations of the WKB behavior of the solutions in the deep subhorizon regime. For perturbative modes above a certain threshold scale $k/\mathcal{H}_{\rm c}$, we observe an enhancement in the transfer functions for the density perturbations. We can attribute this to the period when $c_{s\rm,D}^{2}$ exhibits a dip around the percolation time, as seen in the bottom row of Fig. 3. This behavior is similar to the main results of [Schmid:1998mx], where the vanishing sound speed of the plasma leads to an enhancement in the density perturbations. For benchmarks BP-2 and BP-4, the width of the dip is $\mathcal{H}_{\rm c}(\eta_{f}-\eta_{i})\sim\mathcal{O}\left(10^{-2}\right)$ and $\mathcal{O}\left(10^{-3}\right)$, respectively. Correspondingly, it is expected that those modes with wavenumbers of around $k/\mathcal{H}_{\rm c}\simeq\frac{1}{\mathcal{H}_{\rm c}(\eta_{f}-\eta_{i})}\sim 100$ and $1000$, should have enough resolution to resolve the dip feature in the temporal evolution of the sound speed.
In the limit where $k/\mathcal{H}_{\rm c}\gg 1$, we can understand the growth of the transfer functions by first noting that the density and velocity perturbations at the percolation time can be obtained by using Eq. (4.85), setting $u=u_{*}$. We have
$\left(\begin{matrix}\delta_{*}\\ (k/\mathcal{H}_{\rm c})\hat{\psi}_{*}\end{matrix}\right)=M(k/\mathcal{H}_{\rm c},u_{*})M^{-1}(k/\mathcal{H}_{\rm c},u_{i})T(k)\left(\begin{matrix}\delta_{\rm c}\\ \sqrt{3}\hat{\psi}_{\rm c}\end{matrix}\right)\,.$ (4.93)
Note that
$u_{i}$ $\equiv(-d_{1}-d_{2}\tilde{\eta}_{i})(-d_{2})^{-2/3}=-\left[\frac{k\left(\tilde{\eta}_{*}-\tilde{\eta}_{i}\right)}{\sqrt{3}\mathcal{H}_{\rm c}}\right]^{2/3}\frac{1}{\left(1+3|c_{s\rm,D*}^{2}|\right)^{2/3}}<0\,$, (4.94)
$u_{*}$ $\equiv(-d_{1}-d_{2}\tilde{\eta}_{*})(-d_{2})^{-2/3}=\left[\frac{k\left(\tilde{\eta}_{*}-\tilde{\eta}_{i}\right)}{\sqrt{3}\mathcal{H}_{\rm c}}\right]^{2/3}\frac{3|c_{s\rm,D*}^{2}|}{\left(1+3|c_{s\rm,D*}^{2}|\right)^{2/3}}>0\,.$ (4.95)
In most FOPT scenarios relevant to this work, especially the ones we have featured in Table 1, we have $|c_{s\rm,D*}^{2}|\gg 1$, which may simplify our asymptotic expressions. We then use the asymptotic expansions for the Airy functions and their derivatives, listed in Table 3.
$u>0,u\gg 1$ $u<0,|u|\gg 1$
AiAi $\frac{u^{-1/4}}{2\sqrt{\pi}}\exp\left(-\frac{2u^{3/2}}{3}\right)$ $\frac{|u|^{-1/4}}{\sqrt{\pi}}\cos\left(\frac{2|u|^{3/2}}{3}-\frac{\pi}{4}\right)$
BiBi $\frac{u^{-1/4}}{\sqrt{\pi}}\exp\left(\frac{2u^{3/2}}{3}\right)$ $\frac{|u|^{-1/4}}{\sqrt{\pi}}\sin\left(\frac{2|u|^{3/2}}{3}-\frac{\pi}{4}\right)$
$Ai^{\prime}$ $\frac{u^{1/4}}{2\sqrt{\pi}}\exp\left(-\frac{2u^{3/2}}{3}\right)$ $\frac{|u|^{1/4}}{\sqrt{\pi}}\sin\left(\frac{2|u|^{3/2}}{3}-\frac{\pi}{4}\right)$
$Bi^{\prime}$ $\frac{u^{1/4}}{\sqrt{\pi}}\exp\left(\frac{2u^{3/2}}{3}\right)$ $\frac{|u|^{1/4}}{\sqrt{\pi}}\cos\left(\frac{2|u|^{3/2}}{3}-\frac{\pi}{4}\right)$
Table 3: Asymptotic forms of Airy functions and their derivatives.
We define
@@
$\Delta\chi_{*}\equiv\frac{k}{\mathcal{H}_{\rm c}}|c_{s\rm,D*}^{2}|^{1/2}\Delta\tilde{\eta}_{*}\,,\quad\Delta\tilde{\eta}_{*}\equiv\left(\tilde{\eta}_{*}-\tilde{\eta}_{i}\right)\,$, (4.96)
which is analogous to the argument of the exponential term in the WKB wavefunction, for a particle in the classically forbidden region. One can show that, in the limit $|c_{s\rm,D*}^{2}|\gg 1$, we have
$u_{*}\approx(\Delta\chi_{*})^{2/3}\,,\quad-u_{i}\approx\frac{(\Delta\chi_{*})^{2/3}}{3|c_{s\rm,D*}^{2}|}\,,\quad\left(-d_{2}\right)^{1/3}\approx\frac{(\Delta\chi_{*})^{2/3}}{\Delta\tilde{\eta}_{*}}\,$. (4.97)
In terms of the phase acquired by the perturbations up to $\tilde{\eta}_{i}$,
$\chi_{i}\equiv\frac{k}{\sqrt{3}\mathcal{H}_{\rm c}}\left(1+\tilde{\eta}_{i}\right)\,$, (4.98)
the density and velocity perturbations at percolation are approximately,
@@
$\frac{\delta_{*}}{\mathcal{R}_{k}}$ $\approx\exp\left(\frac{2\Delta\chi_{*}}{3}\right)\Bigg[-\frac{4}{3^{1/4}|c_{s\rm,D*}^{2}|^{1/4}}\left(1+\frac{5}{48\Delta\chi_{*}}\right)\cos\left(\chi_{i}+\frac{\pi}{4}+\frac{2\Delta\chi_{*}}{9\sqrt{3}|c_{s\rm,D*}^{2}|^{3/2}}\right)$
$+\mathcal{O}\left(|c_{s\rm,D*}^{2}|^{5/4}/\Delta\chi_{*}\right)\Bigg]$
$+\exp\left(-\frac{2\Delta\chi_{*}}{3}\right)\left[-\frac{2}{3^{1/4}|c_{s\rm,D*}^{2}|^{1/4}}\sin\left(\chi_{i}+\frac{\pi}{4}+\frac{2\Delta\chi_{*}}{9\sqrt{3}|c_{s\rm,D*}^{2}|^{3/2}}\right)+\mathcal{O}\left(|c_{s\rm,D*}^{2}|^{5/4}/\Delta\chi_{*}\right)\right]\,$,
@@
$\frac{k}{\mathcal{H}_{\rm c}}\frac{\hat{\psi}_{*}}{\mathcal{R}_{k}}$ $\approx\exp\left(\frac{2\Delta\chi_{*}}{3}\right)\Bigg[-\frac{4\Delta\chi_{*}}{3^{1/4}|c_{s\rm,D*}^{2}|^{1/4}\Delta\tilde{\eta}_{*}}\left(1-\frac{7}{48\Delta\chi_{*}}\right)\cos\left(\chi_{i}+\frac{\pi}{4}+\frac{2\Delta\chi_{*}}{9\sqrt{3}|c_{s\rm,D*}^{2}|^{3/2}}\right)$
$+\mathcal{O}\left(|c_{s\rm,D*}^{2}|^{5/4}/\Delta\chi_{*}\right)\Bigg]$
$+\exp\left(-\frac{2\Delta\chi_{*}}{3}\right)\left[\frac{2\Delta\chi_{*}}{3^{1/4}|c_{s\rm,D*}^{2}|^{1/4}\Delta\tilde{\eta}_{*}}\sin\left(\chi_{i}+\frac{\pi}{4}+\frac{2\Delta\chi_{*}}{9\sqrt{3}|c_{s\rm,D*}^{2}|^{3/2}}\right)+\mathcal{O}\left(|c_{s\rm,D*}^{2}|^{5/4}/\Delta\chi_{*}\right)\right]\,.$
The leading terms are
$\frac{\delta_{*}}{\mathcal{R}_{k}}$ $\approx-\frac{4}{(3|c_{s\rm,D*}^{2}|)^{1/4}}~\frac{1}{2}\left(2\epsilon_{+}\cos\chi+\epsilon_{-}\sin\chi\right)\,$, (4.101)
$\frac{k}{\mathcal{H}_{\rm c}}\frac{\hat{\psi}_{*}}{\mathcal{R}_{k}}$ $\approx-\frac{4\Delta\chi_{*}}{(3|c_{s\rm,D*}^{2}|)^{1/4}\Delta\tilde{\eta}_{*}}~\frac{1}{2}\left(2\epsilon_{+}\cos\chi-\epsilon_{-}\sin\chi\right)\,$, (4.102)
${\rm where}\ \ \epsilon_{\pm}$ $\equiv\exp(\pm 2\Delta\chi_{*}/3)\,,\quad\chi\equiv\chi_{i}+\frac{\pi}{4}+\frac{2\Delta\chi_{*}}{9\sqrt{3}|c_{s\rm,D*}^{2}|^{3/2}}\,.$ (4.103)
Keeping the term of order $(\Delta\chi_{*})^{0}$ for $\delta_{*}$ (and order $\Delta\chi_{*}$ for $k/\mathcal{H}_{\rm c}\hat{\psi}_{*}$) amounts to taking the limit $k/\mathcal{H}_{\rm c}\gg 1$. The respective amplitudes of $\delta_{*}$ and $\hat{\psi}_{*}$ are
$\frac{|\delta_{*}|}{\mathcal{R}_{k}}\leq\frac{2}{(3|c_{s\rm,D*}^{2}|)^{1/4}}\sqrt{4\epsilon_{+}^{2}+\epsilon_{-}^{2}}\,,\quad\frac{|\hat{\psi}_{*}|}{\mathcal{R}_{k}}\leq\frac{2(3|c_{s\rm,D*}^{2}|)^{1/4}}{\sqrt{3}}\sqrt{4\epsilon_{+}^{2}+\epsilon_{-}^{2}}\,.$ (4.104)
In Fig. 9 we plot the transfer functions for the density and velocity perturbations for BP-2 and BP-4 as solid curves, along with the amplitude of the approximate transfer functions provided in Eq. (4.104) as dashed green curves. We find that the dashed green curve in both panels captures the exponential growth of the density and velocity perturbations for sufficiently large $k/\mathcal{H}_{\rm c}$. On the other hand, for longer wavelength modes that are still in the subhorizon regime, the amplitudes of the density and velocity transfer functions are close to the amplitudes for the pure radiation background, i.e., 4 and $4/\sqrt{3}$, respectively.
We can understand the overall dependence of the transfer functions with the scale of the perturbations by considering the Airy function approximation to the transfer functions, which we plot in Fig. 10. In this approximation, the shape of the transfer functions are dictated by $\tilde{\eta}_{i}$, $\tilde{\eta}_{*}$, and $c_{s\rm,D*}^{2}$. Furthermore, we assume that we are in the regime where $|c_{s\rm,D*}^{2}|\gg 1$. In the case of the density contrast, the deviation from the expected transfer function in the case of a pure radiation background, according to Eq. (4.99), becomes relevant for scales,
$\exp\left(\frac{2\Delta\chi_{*}}{3}\right)\frac{1}{\left(3|c_{s\rm,D*}^{2}|\right)^{1/4}}\simeq 1\Rightarrow\frac{k}{\mathcal{H}_{\rm c}}\gtrsim\frac{3}{8}\frac{\ln\left(3|c_{s\rm,D*}^{2}|\right)}{|c_{s\rm,D*}^{2}|^{1/2}\Delta\tilde{\eta}_{*}}\,$. (4.105)
As for the velocity perturbation, Eq. (4.100) implies that the deviation from the pure radiation background occurs when
$\exp\left(\frac{2\Delta\chi_{*}}{3}\right)\left(3|c_{s\rm,D*}^{2}|\right)^{1/4}\left(1-\frac{7}{48\Delta\chi_{*}}\right)\simeq 1\Rightarrow\frac{k}{\mathcal{H}_{\rm c}}\gtrsim\frac{7}{48|c_{s\rm,D*}^{2}|^{1/2}\Delta\tilde{\eta}_{*}}\,$. (4.106)
Also, the size of the next to leading order term, relative to the leading term is $\sim|c_{s\rm,D*}^{2}|^{3/2}/\Delta\chi_{*}\sim|c_{s\rm,D*}^{2}|/(\Delta\tilde{\eta}_{*}~k/\mathcal{H}_{\rm c})$. Then $k/\mathcal{H}_{\rm c}$ should be larger than $\sim|c_{s\rm,D*}^{2}|/\Delta\tilde{\eta}_{*}$ to keep the next to leading order smaller than the leading term. In Fig. 10, vertical blue/red dashed lines, respectively, mark the range of modes identified by Eqs. (4.105) and (4.106). For comparison, we mark the scale $k/\mathcal{H}_{\rm c}=1/(\tilde{\eta}_{*}-\tilde{\eta}_{i})$ by a dashed dot line, and the scale $k/\mathcal{H}_{\rm c}=|c_{s\rm,D*}^{2}|/(\tilde{\eta}_{*}-\tilde{\eta}_{i})$ by a dotted line.
The approximate treatment of assuming a linear drop in $c_{s\rm,D}^{2}$ allows us to identify the factors that affect the transfer functions, namely: the period $\Delta\tilde{\eta}_{*}$ when the drop in $c_{s\rm,D}^{2}$ occurs; and the sound speed squared $c_{s\rm,D*}^{2}$ at the percolation time. We demonstrate the effect of these quantites on the transfer functions in Fig. 10. In the top row, we changed $\Delta\tilde{\eta}_{*}$, effectively changing the width of the dip in $c_{s\rm,D}^{2}$ (across columns). As expected, reducing the width simply moves the vertical lines that mark the modes that suffer enhancement in the density to the right. In the bottom row, we changed $c_{s\rm,D*}^{2}$ from -20 to -60. Decreasing $c_{s\rm,D*}^{2}$ has a similar effect of moving the locations of the critical Fourier scales as in the case of changing $\Delta\tilde{\eta}_{*}$, but they are shifted to the left.
@@
$\Delta\tilde{\eta}_{*}=10^{-3},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=-20$ $\Delta\tilde{\eta}_{*}=10^{-4},\tilde{\eta}_{*}=0.1,c_{s\rm,D*}^{2}=-20$